tan a- tan b =x and cot b -cot a=y then cot (a-b) = ?

Grade:10

1 Answers

Shailendra Kumar Sharma
6 years ago
cotb-cota=y
1/tanb-1/tana=y
(tana-tanb)/tanatanb=y
tanatanb=x/y
cot(a-b)=(1+tanatanb)/(tana-tanb)
=(1+x/y)/x
=(x+y)/xy